OpenStudy (anonymous):

70 is what % of 280? Please explain the process. :)

OpenStudy (anonymous):

25% because 70/280 = 1/4 = 0.25

OpenStudy (anonymous):

196, Divide 70 by 280. Move the decimal point two places to the right. Attach a percent sign (%)

OpenStudy (anonymous):

So the answer is 25% because...

OpenStudy (anonymous):

 Is this why?